# A Survey on Dynamic Neural Networks: from Computer Vision to Multi-modal Sensor Fusion

### v1

@@ 2.3.2 Token merging
E-ViT (Liang et al., 2022) performs skimming of the tokens in between the Attention and MLP layers by observing the average of the attention weights of multiple heads at specific blocks (3 blocks, spread in the network). First, the top-$k$ most important tokens are selected, and then all the remaining tokens are merged into one, taking into account a weighted average. The attention weights are used as an indicator of the contribution (and thus importance) of each token to the classification. The importance is estimated with respect to the classification token. Long et al. (2023) improves on top of E-ViT by merging tokens, considering both importance and global token diversity. The proposed technique first splits the tokens into two different sets according to the attention output: attentive tokens and inattentive tokens, as seen also in E-ViT. However, instead of merging all the inattentive tokens, it performs a density peak clustering algorithm for the merging of the inattentive tokens, while a custom cosine similarity-based algorithm is used to cluster the top-$k$ most similar tokens in the attentive group.
@@
The method proposed in Xu et al. (2024) tries to solve the challenge of balancing model efficiency and information preservation for efficient ViTs through the use of Graph-based Token Propagation (GTP). In the paper, the authors propose two different metrics to measure the token’s importance: regeneration difficulty and broadcasting ability. The first one represents the ability of a token to be aggregated by other tokens during the self-attention process, while the second score represents the ability of a token to contribute to other tokens in the self-attention computations. The self-attention map is used to compute these scores, so the method can be applied without learning of other tokens or needing to fine-tune the solution. Based on the spatial position (adjacency) and semantic position (cosine similarity) of tokens, a propagation graph of each image is determined for each input, and the nodes with low scores get merged according to the graph direction. Furthermore, in the paper, it is observed that discarding tokens tends to yield a smoother attention map after Softmax activation. To resolve this issue, attention map sparsification is applied as an anti-oversmoothing mechanism.
@@
Table 4: Summary of Token Skimming techniques presented in this survey.
Paper Summary Type Section
Chen et al. (2021b) Combination of methods to induce token sparsity Hierarchical dropping 2.3.1
Pan et al. (2021) Train per block in a curriculum-learning manner with RL Hierarchical dropping 2.3.1
Rao et al. (2021) Pruning of redundant tokens progressively and dynamically Hierarchical dropping 2.3.1
Fayyaz et al. (2022) Selection based on attention weights of the classification token Hierarchical dropping 2.3.1
Liang et al. (2022) Merges according to the attention weights of multiple heads at specific blocks Merging 2.3.2
Li et al. (2022) Tokens selection based on the weights of the attention block Hierarchical dropping 2.3.1
Mersha (2022) At each block, an function evaluates which tokens should attend it Per-block dropping 2.3.1
Xu et al. (2022) The least important tokens are summarized by a representative token Merging 2.3.2
Yin et al. (2022) Halt tokens by accumulative importance, with bias target exit depth. Hierarchical dropping 2.3.1
Bolya et al. (2023) Gradually combines similar tokens with a custom matching algorithm Merging 2.3.2
Bolya and Hoffman (2023) Applies Bolya et al. (2023) algorithm to diffusion models Merging 2.3.2
Chen et al. (2023) Token dropout for video task recognition Hierarchical dropping 2.3.3
Chen et al. (2023) Introduce activation sparsity for Swin-based models Hierarchical dropping 2.3.3
Havtorn et al. (2023) Selects the optimal token scale for every image region - 2.3.3
Jiang et al. (2023) Token merging based on the similarity of pixel patches Merging 2.3.2
Lei (2023) Per block skip according to the output of inferred score Per-block dropping 2.3.1
Long et al. (2023) Merge tokens according to importance and global token diversity Merging 2.3.2
Lu et al. (2023) Token reduction for ViTs semantic segmentation - 2.3.3
Seol et al. (2023) Skimming based on weighted average according to a importance score of the token Merging 2.3.2
Wei et al. (2023) Scoring method based on Gumbel-Softmax for merging tokens Merging 2.3.2
Devoto et al. (2024) Token skimming for variable latency and bandwidth constraints in communication channels Hierarchical dropping 2.3.3
Kim et al. (2024) Aggregation based on both token pruning and token merging Dropping & merging 2.3.3
Liu et al. (2024) Token Skimming for object detection and semantic segmentation Hierarchical dropping 2.3.3
Peng et al. (2024) Token dropout for Event cameras Hierarchical dropping 2.3.3
Xu et al. (2024) Token merging for Object Tracking Merging 2.3.3
Xu et al. (2024) Uses a Graph-based Token Propagation (GTP) as a merging policy Merging 2.3.2
Xu et al. (2024) Per block skip based on the attention score Per-block dropping 2.3.1
@@ 2.3.3 Other methods and applications
There are numerous other token skimming approaches that do not fall within the token dropping or token merging categories. For example, Chen et al. (2023) is a model that revisits the activation sparsity for Swin-based models (Liu et al., 2021). It introduces a sparsity-aware adaptation and applies an evolutionary search to efficiently find the optimal sparsity for each layer configuration within the vast search space. The importance of each window is given by its L2 activation magnitude. The windows with the highest score are gathered, then passed through the transformer block, and finally, the output is scattered back with the unprocessed windows. Different sparsity levels are applied at different layers due to the principle that early layers need to see more tokens than the later layers. In order to find the optimal sparsity for each layer, the network is first trained to perform at any random sparsity level, then an evolution algorithm is applied to find the optimal sparsity for each layer, and finally, the model is fine-tuned with the identified optimal sparsity configuration.
@@
Chen et al. (2023) proposes a spatiotemporal token dropout to improve the efficiency of video task recognition, preserving keyframes in a video for context and the important motion parts for all other frames. Two pruning stages are performed sequentially: first, for all the non-keyframe tokens, the importance score is calculated with the relative keyframe, and only the top-$k$ tokens are kept. Subsequently, an approach similar to E-ViT (Liang et al., 2022) is used to further skim irrelevant tokens to the prediction.
More recently, Liu et al. (2024) applies the idea of per-block Token Skimming to the tasks of object detection and semantic segmentation. In this case, the selection is performed through sampling from a distribution noised with the Gumbel-Softmax trick to make the decision differentiable. Peng et al. (2024) introduces an hierarchical dropout solution for object detection in scenes recorded by event cameras.
Xu et al. (2024) proposes a token fusion method for the task of object tracking. The input of the encoder is composed of search tokens and exemplar tokens, where the search token represents the object that needs to be tracked and the exemplar token can contain either tracking object tokens, which are useful, or background tokens, which get fused together in order to reduce the overall number of tokens. The selection of the tokens is based on the attention weights.
@@ 3 Dynamic Sensor Fusion in the Context of Computer Vision
Analyzing the number of publications, there is clearly a research gap when compared to other applications of Adaptive Networks. On the other end, we can argue that Dynamic Neural Networks are very beneficial in the context of Sensor Fusion for several reasons. First, they serve efficiency purposes, and usually, Sensor Fusion applications run on low-computing devices or target real-time applications. Introducing some sort of elasticity in elaborating the different inputs allows for reduced computations. It enables to remove redundant and uninformative inputs and relates the use of certain sensors only to cases where the conditions benefit the overall prediction. In this sense, the benefit is derived from a more controlled (and tunable) trade-off between achieving a desired accuracy and constraining the computation resources. Furthermore, dynamically excluding certain sensors can help reduce the noise generated by the said sensors, in cases where the working conditions are suboptimal or a sensor failure is present. In these cases, a dynamic model can perform an improved selection of the optimal subset of sensors, taking into account also the environmental conditions (Chen et al., 2019, Malawade et al., 2022). Finally, these methods are able to identify and prioritize the elaboration of critical information by design. This is particularly useful in contexts where computation is a limited resource and the system needs to be responsive. We can imagine a setting where information can be elaborated either on-device or on the cloud. The introduced dynamicity in this case can help to better differentiate between information that is critical and needs to be computed on-premises for real-time and all-time responsiveness, and the information that can be sent out of the device to improve the overall model accuracy at the cost of latency or possibly information loss. In the following paragraphs, we present the application of Adaptive Sensor Fusion techniques for object detection, semantic segmentation, odometry, steering prediction and finally we introduce two different use cases of Sensor Fusion in the context of stress detection and power grid attack detection.
@@
Malawade et al. (2022) presents HydraFusion, a framework that applies dynamic late fusion (after Faster-RCNN sensor processing) in a context-aware manner to perform object detection. It selects a fixed amount of $k$ sensors out of $n$, and uses a novel attention-based module to determine which are the most important sensors. The execution time of the model is on par with any late fusion method due to the fact that all inputs are processed but only $k$ of them are selected. This selection brings improvements in accuracy due to the induced robustness of the architecture in cherry-picking only a subset of inputs. It is also noticeable that in the case where $k=1$ (only one sensor selected among all), the ablation study shows that a statistical method based on domain knowledge (based on the weather conditions) is sufficient while increasing $k$ (multiple sensor selected and fused) deep learning-based selection performs better.
Similarly, EcoFusion (Malawade et al., 2022) proposes an energy-aware Sensor Fusion approach which uses the environmental context to adapt the fusion method to avoid the negative impact of the fusion process while reducing the energy consumption. The method is able to dynamically switch between different sensor combinations. The optimization is formulated as a joint minimization problem between energy consumption and model loss. A hyperparameter allows to represent the maximum deviation from the maximum performing configuration, exploring more efficient configurations. The workflow goes as follows: sensor measurements are passed through modality-specific stem models, which produce an initial set of features. A gate model then estimates the loss based on the set of features and the possible fusion configurations. The best optimization approaches are then selected, running the select branches for object detection, the outputs of which are then fused in a late fusion block. Four different gating approaches have been tested: knowledge-based, MLP-based, attention-based, and loss-based. This flexibility allows the model to dynamically choose between early fusion, late fusion, no fusion, or a combination of the three.
@@
More recently, Zhang et al. (2024) introduces the technique of token merging in the case of missing modalities for the task of brain tumor segmentation. The core idea is to reduce the model size by extracting and merging accessible modalities into more compact token sequences. The architecture incorporates two main blocks: a unimodal token merging block focuses on information extraction from individual modalities by compressing spatially redundant tokens, and a multi-modal token merging block, responsible for the fusion of the tokens from the different modalities to enhance global representation.
@@
Chen et al. (2019) propose an end-to-end Sensor Fusion framework for monocular visual-inertial odometry that fuses selectively monocular images and inertial measurements (IMU). This allows to estimate the trajectory taking into account missing and corrupted data. The model proposes two fusion methods: a deterministic soft fusion and a hard fusion method based on the Gumbel-Softmax resampling. In particular, this last one allows to generate a binary mask that either propagates the feature or blocks it according to the environment and its reliability.
@@
CARMA (Zhang et al., 2023) is a framework proposing a fusion approach that uses the context to dynamically reconfigure the computation flow on a field-programmable gate array (FPGA) at runtime. This is performed through clock-gating unused sensors and model sub-components, with the aim of reducing the energy consumption. The architecture allows an optimization of the whole system, including sensors, model architecture, and hardware platform. A context identification process is enabled intermittently. The advantage of such a solution is to have global control over the overall power consumption of the whole system and not only the model part.
@@
Paper Summary Task
Mees et al. (2016) Convolutional MoE for RGB, depth and motion Object detection
Valada et al. (2017) Convolutional MoE to dynamically fuse different modalities Semantic segmentation
Blum et al. (2018) Late fusion approach for semantic segmentation from the output of separately trained experts Semantic segmentation
John et al. (2018) Road scenes and driving patterns based fusion for steering angle estimation Steering prediction
Chen et al. (2019) Selective fusion of images and IMU Odometry
Chen et al. (2022) Dynamic offload of sensor process to edge computing units Object detection
Malawade et al. (2022) Adds the environmental context to Dynamic Sensor Fusion Object detection
Malawade et al. (2022) Sensor fusion selection to perform object detection Object detection
Cao et al. (2023) Multi-modal mixture of local-to-global experts Object detection
Morra et al. (2023) Local optimal expert selection for multicamera visual odometry Odometry
Rashid et al. (2023) context-aware Sensor Fusion for stress detection on embedded devices Stress detection
Zhang et al. (2023) Fusion approach to dynamically reconfigure FPGA at runtime Any
Mortlock and Al Faruque (2024) Fusion for power grid state estimation to create robustness towards attacks Regression
Zhang et al. (2024) Token merging for brain tumor segmentation Semantic segmentation

### v3

@@ 2.3.2 Token merging
E-ViT (Liang et al., 2022b) performs skimming of the tokens in between the Attention and MLP layers by observing the average of the attention weights of multiple heads at specific blocks (3 blocks, spread in the network). First, the top-$k$ most important tokens are selected, and then all the remaining tokens are merged into one, taking into account a weighted average. The attention weights are used as an indicator of the contribution (and thus importance) of each token to the classification. The importance is estimated with respect to the classification token. Long et al. (2023) improves on top of E-ViT by merging tokens, considering both importance and global token diversity. The proposed technique first splits the tokens into two different sets according to the attention output: attentive tokens and inattentive tokens, as seen also in E-ViT. However, instead of merging all the inattentive tokens, it performs a density peak clustering algorithm for the merging of the inattentive tokens, while a custom cosine similarity-based algorithm is used to cluster the top-$k$ most similar tokens in the attentive group.
@@
The method proposed in Xu et al. (2024c) tries to solve the challenge of balancing model efficiency and information preservation for efficient ViTs through the use of Graph-based Token Propagation (GTP). In the paper, the authors propose two different metrics to measure the token’s importance: regeneration difficulty and broadcasting ability. The first one represents the ability of a token to be aggregated by other tokens during the self-attention process, while the second score represents the ability of a token to contribute to other tokens in the self-attention computations. The self-attention map is used to compute these scores, so the method can be applied without learning of other tokens or needing to fine-tune the solution. Based on the spatial position (adjacency) and semantic position (cosine similarity) of tokens, a propagation graph of each image is determined for each input, and the nodes with low scores get merged according to the graph direction. Furthermore, in the paper, it is observed that discarding tokens tends to yield a smoother attention map after Softmax activation. To resolve this issue, attention map sparsification is applied as an anti-oversmoothing mechanism.
Yuan et al. (2024) introduces a token merging system based on a self-attentive method. The introduced block creates a score matrix that compares each input token with each other, subsequently allowing the merging of tokens through a weighted sum. An inflation block is also introduced to reconstruct the tokens after the transformer block.
@@
Table 4: Summary of Token Skimming techniques presented in this survey (1/2).
Paper Summary Type Section Chen et al. (2021b) Combination of methods to induce token sparsity Hierarchical dropping 2.3.1 Pan et al. (2021) Train per block in a curriculum-learning manner with RL Hierarchical dropping 2.3.1 Rao et al. (2021) Pruning of redundant tokens progressively and dynamically Hierarchical dropping 2.3.1 Fayyaz et al. (2022) Selection based on attention weights of the classification token Hierarchical dropping 2.3.1 Liang et al. (2022b) Merges according to the attention weights of multiple heads at specific blocks Merging 2.3.2 Li et al. (2022) Tokens selection based on the weights of the attention block Hierarchical dropping 2.3.1 Mersha (2022) At each block, an function evaluates which tokens should attend it Per-block dropping 2.3.1 Xu et al. (2022) The least important tokens are summarized by a representative token Merging 2.3.2 Yin et al. (2022) Halt tokens by accumulative importance, with bias target exit depth. Hierarchical dropping 2.3.1 Bolya et al. (2023) Gradually combines similar tokens with a custom matching algorithm Merging 2.3.2 Bolya and Hoffman (2023) Applies Bolya et al. (2023) algorithm to diffusion models Merging 2.3.2 Chen et al. (2023a) Token dropout for video task recognition Hierarchical dropping 2.3.3 Chen et al. (2023b) Introduce activation sparsity for Swin-based models Hierarchical dropping 2.3.3 Havtorn et al. (2023) Selects the optimal token scale for every image region - 2.3.3 Jiang et al. (2023) Token merging based on the similarity of pixel patches Merging 2.3.2 Lei et al. Per block skip according to the output of inferred score Per-block dropping 2.3.1 Long et al. (2023) Merge tokens according to importance and global token diversity Merging 2.3.2 Lu et al. (2023) Token reduction for ViTs semantic segmentation - 2.3.3 Seol et al. (2023) Skimming based on weighted average according to a importance score of the token Merging 2.3.2 Wei et al. (2023) Scoring method based on Gumbel-Softmax for merging tokens Merging 2.3.2 Devoto et al. (2024) Token skimming for variable latency and bandwidth constraints in communication channels Hierarchical dropping 2.3.3
Table 4: Summary of Token Skimming techniques presented in this survey(2/2).
Paper Summary Type Section Kim et al. (2024) Aggregation based on both token pruning and token merging Dropping & merging 2.3.3 Liu et al. (2024b) Token Skimming for object detection and semantic segmentation Hierarchical dropping 2.3.3 Peng et al. (2024) Token dropout for Event cameras Hierarchical dropping 2.3.3 Xu et al. (2024a) Token merging for Object Tracking Merging 2.3.3 Xu et al. (2024c) Uses a Graph-based Token Propagation (GTP) as a merging policy Merging 2.3.2 Xu et al. (2024b) Per block skip based on the attention score Per-block dropping 2.3.1 Yuan et al. (2024) Token merging system base on self-attentive method Merging 2.3.2 Cheng et al. (2025) Controls the token sparsity through the use of the importance distribution for object detection Merging 2.3.3 Fang et al. (2025) Merge redundant features in the context of diffusion models Merging 2.3.3 Lu et al. (2025) Token merging for diffusion transformers with GPU friendly operations Merging 2.3.3
@@ 2.3.3 Other methods and applications
There are numerous other token skimming approaches that do not fall within the token dropping or token merging categories. For example, Chen et al. (2023b) is a model that revisits the activation sparsity for Swin-based models (Liu et al., 2021a). It introduces a sparsity-aware adaptation and applies an evolutionary search to efficiently find the optimal sparsity for each layer configuration within the vast search space. The importance of each window is given by its L2 activation magnitude. The windows with the highest score are gathered, then passed through the transformer block, and finally, the output is scattered back with the unprocessed windows. Different sparsity levels are applied at different layers due to the principle that early layers need to see more tokens than the later layers. In order to find the optimal sparsity for each layer, the network is first trained to perform at any random sparsity level, then an evolution algorithm is applied to find the optimal sparsity for each layer, and finally, the model is fine-tuned with the identified optimal sparsity configuration.
@@
Chen et al. (2023a) proposes a spatiotemporal token dropout to improve the efficiency of video task recognition, preserving keyframes in a video for context and the important motion parts for all other frames. Two pruning stages are performed sequentially: first, for all the non-keyframe tokens, the importance score is calculated with the relative keyframe, and only the top-$k$ tokens are kept. Subsequently, an approach similar to E-ViT (Liang et al., 2022b) is used to further skim irrelevant tokens to the prediction.
More recently, Liu et al. (2024b) applies the idea of per-block Token Dropping to the tasks of object detection and semantic segmentation. In this case, the selection is performed through sampling from a distribution noised with the Gumbel-Softmax trick to make the decision differentiable. Cheng et al. (2025) controls the token sparsity through the use of the importance distribution. This allows the solution to preserve both spatial details and the global receptive field when used with a DETR head Carion et al. (2020). Peng et al. (2024) introduces an hierarchical dropout solution for object detection in scenes recorded by event cameras.
Xu et al. (2024a) proposes a token fusion method for the task of object tracking. The input of the encoder is composed of search tokens and exemplar tokens, where the search token represents the object that needs to be tracked and the exemplar token can contain either tracking object tokens, which are useful, or background tokens, which get fused together in order to reduce the overall number of tokens. The selection of the tokens is based on the attention weights.
Fang et al. (2025) investigates how to prune feature redundancies in the context of diffusion models, introducing a structure-then-detail token merging approach. Lu et al. (2025) focuses on token merging for diffusion transformers, taking into account only GPU-friendly operations.
@@ 3 Dynamic Sensor Fusion in the Context of Computer Vision
Analyzing the number of publications, there is clearly a research gap when compared to other applications of Adaptive Networks. On the other end, we can argue that Dynamic Neural Networks are very beneficial in the context of Sensor Fusion for several reasons. First, they serve efficiency purposes, and usually, Sensor Fusion applications run on low-computing devices or target real-time applications. Introducing some sort of elasticity in elaborating the different inputs allows for reduced computations. It enables to remove redundant and uninformative inputs and relates the use of certain sensors only to cases where the conditions benefit the overall prediction. In this sense, the benefit is derived from a more controlled (and tunable) trade-off between achieving a desired accuracy and constraining the computation resources. Furthermore, dynamically excluding certain sensors can help reduce the noise generated by the said sensors, in cases where the working conditions are suboptimal or a sensor failure is present. In these cases, a dynamic model can perform an improved selection of the optimal subset of sensors, taking into account also the environmental conditions (Chen et al., 2019a; Malawade et al., 2022b). Finally, these methods are able to identify and prioritize the elaboration of critical information by design. This is particularly useful in contexts where computation is a limited resource and the system needs to be responsive. We can imagine a setting where information can be elaborated either on-device or on the cloud. The introduced dynamicity in this case can help to better differentiate between information that is critical and needs to be computed on-premises for real-time and all-time responsiveness, and the information that can be sent out of the device to improve the overall model accuracy at the cost of latency or possibly information loss. In the following paragraphs, we present the application of Adaptive Sensor Fusion techniques for object detection, semantic segmentation, odometry, steering prediction and finally we introduce three different use cases of Sensor Fusion in the context of stress detection, power grid attack detection and multi-modal medical diagnosis.
@@
Malawade et al. (2022a) presents HydraFusion, a framework that applies dynamic late fusion (after Faster-RCNN sensor processing) in a context-aware manner to perform object detection. It selects a fixed amount of $k$ sensors out of $n$, and uses a novel attention-based module to determine which are the most important sensors. The execution time of the model is on par with any late fusion method due to the fact that all inputs are processed but only $k$ of them are selected. This selection brings improvements in accuracy due to the induced robustness of the architecture in cherry-picking only a subset of inputs. It is also noticeable that in the case where $k=1$ (only one sensor selected among all), the ablation study shows that a statistical method based on domain knowledge (based on the weather conditions) is sufficient while increasing $k$ (multiple sensor selected and fused) deep learning-based selection performs better.
Similarly, EcoFusion (Malawade et al., 2022b) proposes an energy-aware Sensor Fusion approach which uses the environmental context to adapt the fusion method to avoid the negative impact of the fusion process while reducing the energy consumption. The method is able to dynamically switch between different sensor combinations. The optimization is formulated as a joint minimization problem between energy consumption and model loss. A hyperparameter allows to represent the maximum deviation from the maximum performing configuration, exploring more efficient configurations. The workflow goes as follows: sensor measurements are passed through modality-specific stem models, which produce an initial set of features. A gate model then estimates the loss based on the set of features and the possible fusion configurations. The best optimization approaches are then selected, running the select branches for object detection, the outputs of which are then fused in a late fusion block. Four different gating approaches have been tested: knowledge-based, MLP-based, attention-based, and loss-based. This flexibility allows the model to dynamically choose between early fusion, late fusion, no fusion, or a combination of the three.
@@
More recently, Zhang et al. (2024b) introduces the technique of token merging in the case of missing modalities for the task of brain tumor segmentation. The core idea is to reduce the model size by extracting and merging accessible modalities into more compact token sequences. The architecture incorporates two main blocks: a unimodal token merging block focuses on information extraction from individual modalities by compressing spatially redundant tokens, and a multi-modal token merging block, responsible for the fusion of the tokens from the different modalities to enhance global representation.
@@
Chen et al. (2019a) propose an end-to-end Sensor Fusion framework for monocular visual-inertial odometry that fuses selectively monocular images and inertial measurements (IMU). This allows to estimate the trajectory taking into account missing and corrupted data. The model proposes two fusion methods: a deterministic soft fusion and a hard fusion method based on the Gumbel-Softmax resampling. In particular, this last one allows to generate a binary mask that either propagates the feature or blocks it according to the environment and its reliability.
@@
CARMA (Zhang et al., 2023b) is a framework proposing a fusion approach that uses the context to dynamically reconfigure the computation flow on a field-programmable gate array (FPGA) at runtime. This is performed through clock-gating unused sensors and model sub-components, with the aim of reducing the energy consumption. The architecture allows an optimization of the whole system, including sensors, model architecture, and hardware platform. A context identification process is enabled intermittently. The advantage of such a solution is to have global control over the overall power consumption of the whole system and not only the model part.
@@
AdaCoMed (Chen et al., 2025) presents a framework that combines different sized models for medical diagnosis. The paper presents a mix of Vision and Language models through the use of a MoE architecture. It combines features extracted from multiple single-modal medical large models in combination with multi-modal small models, through a co-learning mechanism which properly weights the strengths of each type of architecture.
@@
Paper Summary Sensors Method Task Mees et al. (2016) Convolutional MoE for RGB, depth and motion RGB, Depth Mixture-of-Experts Object detection Valada et al. (2017) Convolutional MoE to dynamically fuse different modalities RGB, Depth, EVI Mixture-of-Experts Semantic segmentation Blum et al. (2018) Late fusion approach for semantic segmentation from the output of separately trained experts RGB, Depth Mixture-of-Experts Semantic segmentation John et al. (2018) Road scenes and driving patterns based fusion for steering angle estimation RGB Mixture-of-Experts Steering prediction Chen et al. (2019a) Selective fusion of images and IMU RGB, IMU Gumbel-Softmax fusion Odometry Chen et al. (2022) Dynamic offload of sensor process to edge computing units RGB Stereo, LiDAR, Radar Early Exits Object detection Malawade et al. (2022b) Adds the environmental context to Dynamic Sensor Fusion RGB, LiDAR, Radar Dynamic Architecture Object detection Malawade et al. (2022a) Sensor fusion selection to perform object detection RGB, LiDAR, Radar Dynamic Architecture Object detection Cao et al. (2023) Multi-modal mixture of local-to-global experts RGB, IR Mixture-of-Experts Object detection Morra et al. (2023) Local optimal expert selection for multicamera visual odometry RGB Mixture-of-Experts Odometry Rashid et al. (2023) Context-aware Sensor Fusion for stress detection on embedded devices Wearable Sensors Mixture-of-Experts Stress detection Zhang et al. (2023b) Fusion approach to dynamically reconfigure FPGA at runtime Any Dynamic Architecture Any Mortlock and Al Faruque (2024) Fusion for power grid state estimation to create robustness towards attacks Grid Power Sensors Dynamic Architecture Regression Zhang et al. (2024b) Token merging for brain tumor segmentation FLAIR, T1ce, T1, and T2 Token merging Semantic segmentation